Suppose Latte Larrys successfully eliminates their rival Mocha Joes so that Latte Larrys is now a coffee monopolist. Let L
Answer #1

99 = 29 P(q) = 20- 049 1 Monopalys profit - MR = MC - TR = P(q) (2) (20-0.49 TR = 202-0.49² TMR = 20-0.89 ] TTC = 29. I MC =